# Multi-Objective min-max Online Convex Optimization

### v1

@@ Abstract
In online convex optimization (OCO), a single loss function sequence is revealed over a time horizon of $T$, and an online algorithm has to choose its action at time $t$, before the loss function at time $t$ is revealed. The goal of the online algorithm is to incur minimal penalty (called regret) compared to a static optimal action made by an optimal offline algorithm knowing all functions of the sequence in advance.
In this paper, we broaden the horizon of OCO, and consider multi-objective OCO, where there are $K$ distinct loss function sequences, and an algorithm has to choose its action at time $t$, before the $K$ loss functions at time $t$ are revealed. To capture the tradeoff between tracking the $K$ different sequences, we consider the min-max regret, where the benchmark (optimal offline algorithm) takes a static action across all time slots that minimizes the maximum of the total loss (summed across time slots) incurred by each of the $K$ sequences. An online algorithm is allowed to change its action across time slots, and its min-max regret is defined as the difference between its min-max cost and that of the benchmark. The min-max regret is a stringent performance measure and an algorithm with small regret needs to ‘track’ all loss function sequences closely at all times.
We consider this min-max regret in the i.i.d. input setting where all loss functions are i.i.d. generated from an unknown distribution. For the i.i.d. model we propose a simple algorithm that combines the well-known Hedge and online gradient descent (OGD) and show via a remarkably simple proof that its expected min-max regret is $O(\sqrt{T\log K})$.
@@ 1 Introduction
Online convex optimization (OCO) Hazan (2019) is a basic learning problem that models a wide variety of applications such as spam filtering Hazan (2019), capacity provisioning in cloud networks Lin et al. (2012) etc., where at time $t$, an online algorithm $\mathcal{A}$ has to choose an action $x_{t}\in\mathcal{X}$, after which the adversary reveals the convex objective/loss function $f_{t}$ on which $x_{t}$ is evaluated. As a benchmark, the offline optimal algorithm is considered, that knows the entire set of $f_{t}$, $t=1,\dots,T$, however, can only choose a single action across all times. To quantify the performance of online algorithms, the metric of static regret $R^{\text{s}ingle}_{T}$ is studied, that is defined as follows, where $x_{t}$’s are the actions taken by $\mathcal{A}$,
$R^{\text{s}ingle}_{\mathcal{A}}(T)=\sup_{f_{t}}\left\{\sum_{t=1}^{T}f_{t}(x_{t})-\min_{x}\sum_{t=1}^{T}f_{t}(x)\right\}.$ (1)
In a long line of work, comprehensively summarized in Hazan (2019), tight guarantees (both upper and lower bounds) have been obtained for the OCO problem, and $R^{\text{s}ingle}(T)=\inf_{\mathcal{A}}R^{\text{s}ingle}_{\mathcal{A}}(T)=\Theta(\sqrt{T})$. In particular, online gradient descent (OGD), mirror descent, and follow-the-regularized-leader (FTRL) Hazan (2019) all achieve the optimal static regret performance for OCO.
@@
Thus, in this paper, we consider the robust alternative in the multi-objective case: the $\it min-max$} OCO, where the objective of an algorithm is to minimize the maximum of the total loss (summed across time slots) incurred by each of the multiple objective functions. A compelling example for considering the min-max approach is fair machine learning, where performance is typically evaluated across multiple demographic groups, which naturally induces a multi-objective optimization problem. While Pareto efficiency identifies trade-offs between groups, it may still permit highly unequal solutions, as a classifier can be Pareto optimal even if it performs very poorly on a disadvantaged group. To address this limitation, fairness is often modeled via a min–max objective: maximize the minimum accuracy (or minimize the maximum error) across groups Chen et al. (2020); Chouldechova and Roth (2020)).
The min-max approach is also well suited for modern networked systems—such as data centers, wireless edge networks, and shared cloud infrastructures—where online decisions must balance performance across multiple service classes, users, or flows whose objectives (throughput, latency, energy) often conflict. The min–max regret criterion provides a principled way to ensure worst-case fairness by minimizing the largest cumulative loss among competing entities, which is crucial in fair resource allocation Sinha et al. (2023), fair allocation in clouds Hao et al. (2016), fair assignment of reusable resources Liu and Hajiesmaili (2025) congestion-controlled transport Kouvelis and Yu (1997); Bertsimas et al. (2011), and dynamic resource provisioning in data centers Urgaonkar et al. (2010). Unlike average-regret formulations that can conceal systematic bias, minimizing min–max regret explicitly guards against persistent under-allocation to any group or flow—an increasingly important property in multi-tenant and QoS-sensitive networks.
The formal min-max OCO problem that we consider is as follows. We consider that there are two convex loss function sequences $f_{t},g_{t},$ that are revealed after $\mathcal{A}$ chooses action $x_{t}$ at time $t$ for $t=1,\dots,T$. The min-max cost of $\mathcal{A}$ is
@@
Two loss function sequences are considered for simplicity of exposition since they capture the basic difficulty. Extension to $K>2$ sequences is straightforward and detailed in Remark 7. As a benchmark, we consider the static offline optimal algorithm $\mathsf{OPT}$ that knows $f_{t},g_{t}$ for $t=1,\dots,T$ at time $t=1$ itself but can only choose a single action $x^{\star}$ across all times such that
$x^{\star}=\arg\min_{x\in\mathcal{X}}\max\left\{\sum_{t=1}^{T}f_{t}(x),\sum_{t=1}^{T}g_{t}(x)\right\},$ (3)
and the optimal cost is
$C_{\mathsf{OPT}}=\max\left\{\sum_{t=1}^{T}f_{t}(x^{\star}),\sum_{t=1}^{T}g_{t}(x^{\star})\right\}.$ (4)
Consequently, we define the min-max static regret of $\mathcal{A}$ with actions $x_{t}$ as
@@
Remark 1 (On the regret definition).
The chosen benchmark in (5) is deliberately stringent. If we considered a weaker benchmark, e.g., minimizing $\min_{x\in X}\sum_{t=1}^{T}\max\{f_{t}(x),g_{t}(x)\}$), then applying standard OGD algorithm on the surrogate function $h_{t}(x)=\max\{f_{t}(x),g_{t}(x)\}$ would yield regret $O(\sqrt{T})$ independent of the number of objectives, underestimating the intrinsic hardness of balancing multiple sequences. In particular:
  • Our definition forces the algorithm to track both sequences globally and simultaneously. This means the online algorithm cannot focus on one objective at the expense of the other, nor can it switch actions adaptively across time as the per-slot benchmark allows.
  • Our benchmark ensures robustness across objectives: the comparator $x^{\star}$ is the best fixed action that minimizes the worst cumulative cost.
  • This definition aligns with fairness and robustness considerations (e.g., in machine learning across demographic groups), where good performance must be guaranteed for all objectives, not merely on average or in expectation.
Thus, while the regret metric in (5) is harder to minimize, it is an appropriate notion when robustness across multiple objectives is essential.
1.1 Related Work
Because of its widespread applications, multi-objective OCO has been an object of immense interest, however, with limited success. For example, Mahdavi et al. (2013) wanted to study the multi-objective OCO and stated the difficulty as “A fundamental difference between single- and multi-objective optimization is that for the latter it is not obvious how to evaluate the optimization quality”. Further, it said “Since it is impossible to simultaneously minimize multiple loss functions and in order to avoid complications caused by handling more than one objective, we choose one function as the objective and try to bound other objectives by appropriate thresholds. This work gave rise to the now popular problem of constrained online convex optimization (COCO), where the problem is transformed into a constrained optimization problem with one objective and one or more constraint functions, that is described next.
1.1.1 COCO
$\textrm{CCV}_{\mathcal{A}}(T)\equiv\sum_{t=1}^{T}\max(g_{t}(x_{t}),0).$
1.1.2 Prior Work on COCO
(A) i.i.d. loss and constraints: COCO with independent and identically distributed (i.i.d.) loss functions $f_{t}$’s and constraint functions $g_{t}$’s was studied in Mahdavi et al. (2013) while its extension to stationary and ergodic processes can be found in Uziel and El-Yaniv (2017), where both derived an algorithm with regret $O(\sqrt{T})$ and CCV $O(T^{3/4})$.
(B) Time-invariant constraints: COCO with time-invariant constraints, i.e., $g_{t}=g,\forall\ t$ (Yuan and Lamperski, 2018; Jenatton et al., 2016; Mahdavi et al., 2012; Yi et al., 2021) has been considered extensively, where functions $g$ are assumed to be known to the algorithm a priori. The best known results in this case give static regret of $O(T^{\max\{c,1-c\}})$ and CCV of $O(T^{(1-c)/2})$ for any $c\in(0,1)$. The algorithm is allowed to take actions that are infeasible at any time to avoid the costly projection step of the vanilla projected OGD algorithm and the main objective was to design an efficient algorithm with a small regret and CCV while avoiding the explicit projection step.
(C) Time-varying constraints: The more difficult question is solving COCO problem when the constraint functions, i.e., $g_{t}$’s, change arbitrarily with time $t$. One popular algorithm for solving COCO considered a Lagrangian function optimization that is updated using the primal and dual variables (Yu et al., 2017; Sun et al., 2017; Yi et al., 2023). Alternatively, (Neely and Yu, 2017) and (Liakopoulos et al., 2019) used the drift-plus-penalty (DPP) framework (Neely, 2010) to solve the COCO, but which needed additional assumption, e.g. the Slater’s condition in (Neely and Yu, 2017) and with weaker form of the feasibility assumption in (Neely and Yu, 2017).
(Guo et al., 2022) obtained bounds similar to (Neely and Yu, 2017) but without assuming Slater’s condition. Very recently, the state of the art guarantees on simultaneous bounds on regret $O(\sqrt{T})$ and CCV $O(\sqrt{T}\log T)$ for COCO were derived in (Sinha and Vaze, 2024) with a very simple algorithm that combines the loss function at time $t$ and the CCV accrued till time $t$ in a single loss function, and then executes the online gradient descent (OGD) algorithm on the single loss function with an adaptive step-size. Most recently (Sinha and Vaze, 2025) showed lower CCV is possible with increased regret compared to (Sinha and Vaze, 2024) while (Vaze and Sinha, 2025) obtained instance dependent CCV bounds which can be $O(1)$ for specific cases while achieving regret of $O(\sqrt{T})$, respectively. The COCO problem has been considered in the dynamic setting as well (Chen and Giannakis, 2018; Cao and Liu, 2018; Vaze, 2022; Liu et al., 2022) where the benchmark $x^{\star}$ in (6) is replaced by $x_{t}^{\star}$ ($x_{t}^{\star}=\arg\min_{x}f_{t}(x)$) that is also allowed to change its actions over time.
@@ 1.2 Global Cost Functions in Experts Setting
A special case of the min-max OCO problem (5) was considered in Even-Dar et al. (2009), where there are $K$ experts, with expert $i$’s loss at time $t$, being $\ell_{t}(i)$. If expert $i$ is selected by an algorithm $\mathcal{A}$ with probability $p_{t}(i)$ at time $t$, then the overall expected loss for expert $i$ is $L_{i}=\sum_{t=1}^{T}p_{t}(i)\ell_{t}(i)$ where $T$ is the time horizon, and the cost of $\mathcal{A}$ is the global cost Even-Dar et al. (2009)
$C_{\mathcal{A}}=\max_{i=1,\dots,K}\{L_{i}\}.$ (7)
The regret of $\mathcal{A}$ is defined as usual with respect to the static optimal offline algorithm that is required to choose a fixed distribution across all time slots while knowing all losses in advance.
The name global refers to the fact that the objective function (7) captures the end-to-end cost function rather than a summation of the instantaneous losses. For this formulation, which we refer to as GlobalExpertsProblem, Even-Dar et al. (2009) proposed an algorithm called MULTI and showed its regret to be $O(\sqrt{T}\log(K))$. GlobalExpertsProblem differs with our min-max OCO formulation, in two fundamental ways, i) the objective inside the $\max$ is a linear function of the actions $p_{t}(i)$, and ii) the action $p_{t}(i)$ taken at time $t$ affects each expert’s loss but only in an average sense since $\sum_{i=1}^{K}p_{t}(i)=1$ while with min-max OCO the action $x_{t}$ chosen at time $t$ affects both $\sum_{t=1}^{T}f_{t}(x_{t})$ and $\sum_{t=1}^{T}g_{t}(x_{t})$ in full. Even though this appears to only be a small difference, it changes everything. As a consequence the result on GlobalExpertsProblem does not apply for min-max OCO even when $f_{t},g_{t}$ are linear. Extensions of Even-Dar et al. (2009) with additive global functions can be found in Azar et al. (2014); Kesselheim and Singla (2020) while for non-additive ones in Rakhlin et al. (2011).
@@ 1.3 Pareto Optimal Multi-Objective OCO
1.4 Per-slot $\it{\it min-max}$} multi-objective optimization
Compared to our ‘global’ min-max objective (4), in Lee et al. (2022), a per-slot $\it min-max$} multi-objective OCO was studied, where in particular, at each round $t=1,\dots,T$: i) an online algorithm chooses $x_{t}\in\mathcal{X}$ (convex compact set), ii) the adversary chooses $y_{t}\in\mathcal{Y}$ (convex compact set), and iii) the online algorithm suffers a vector loss:
$\ell_{t}(x_{t},y_{t})=\big(\ell_{t,1}(x_{t},y_{t}),\dots,\ell_{t,d}(x_{t},y_{t})\big)\in[-1,1]^{d}.$
The goal of an online algorithm is to solve $\min_{\{x_{t}\}}\max_{j\in[d]}\sum_{t=1}^{T}\ell_{t,j}(x_{t},y_{t})$, knowing $\ell_{t}(.,.)$, but not $y_{t}$, and its performance is compared against a per-round local minimax value
$w_{L}^{t}:=\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\max_{j\in[d]}\ell_{t,j}(x,y),$
where the cumulative benchmark is:
@@ 1.4 Per-slot $\it{\it min-max}$} multi-objective optimization
A Hedge-type algorithm Cesa-Bianchi and Lugosi (2006) was shown in Lee et al. (2022) to have $O(\sqrt{T})$ regret.
Defining $\mathcal{Y}=\{e\}$ a singleton, and $d=2$, such that $\ell_{t,1}(x_{t},y_{t})=f_{t}(x_{t})$ and $\ell_{t,1}(x_{t},y_{t})=g_{t}(x_{t})$ for $y_{t}=e$, we recover our model where two convex functions $f_{t},g_{t}$ are revealed at time $t$, and for which the benchmark $W_{L}^{T}$ simplifies to
$W_{L}^{T}=\sum_{t=1}^{T}\min_{x_{t}\in\mathcal{X}}\max\{f_{t}(x_{t}),g_{t}(x_{t})\}).$ (10)
Thus, benchmark (10) is fundamentally different than our considered benchmark (4) since it is additive across time slots, and allows the benchmark to choose different actions across time slots. Even though (10) allows choosing different actions across time slots, surprisingly, we show in Appendix D that $W_{L}^{T}$ (10) can be $3/2$ times more than our benchmark $C_{\mathsf{OPT}}$ (4) and hence algorithms with $O(\sqrt{T})$ regret (9) do not yield a $O(\sqrt{T})$ min-max static regret (5).
In summary, compared to COCO and GlobalExpertsProblem, our considered problem has a wider scope, while our considered metric (5) is more meaningful and stringent compared to (8) and (10).
1.5 Applications
Multi-objective OCO has also been studied from an applications perspective, e.g. Lyu et al. (2019) studied the problem of multi-period multi-objective online ride-matching problem, where the multiple objectives included platform revenue, assigning drivers with higher service quality, minimizing the distance to the customer etc. Lyu et al. (2019) also used the COCO-type reformulation of multi-objective OCO for deriving theoretical guarantees. Other applications of multi-objective OCO include fair resource allocation Sinha et al. (2023), fair allocation in clouds Hao et al. (2016), fair assignment of reusable resources Liu and Hajiesmaili (2025) congestion-controlled transport Kouvelis and Yu (1997); Bertsimas et al. (2011), and dynamic resource provisioning in data centers Urgaonkar et al. (2010), multi-task learning Sener and Koltun (2018), multiple target tracking Milan et al. (2017), portfolio selection Das and Banerjee (2011) and safe reinforcement learning Chow et al. (2018), where decision-makers must make sequential decisions under uncertainty, balancing objectives such as cost, delivery time, and reliability without full knowledge of future demand or disruptions Kouvelis and Yu (1997); Bertsimas et al. (2011), where the min-max criterion naturally makes sense.

### v3

@@ Abstract
In this paper, we broaden the horizon of online convex optimization (OCO), and consider multi-objective OCO, where there are $K$ distinct loss function sequences, and an algorithm has to choose its action at time $t$, before the $K$ loss functions at time $t$ are revealed. To capture the tradeoff between tracking the $K$ different sequences, we consider the min-max regret, where the benchmark (optimal offline algorithm) takes a static action across all time slots that minimizes the maximum of the total loss (summed across time slots) incurred by each of the $K$ sequences. An online algorithm is allowed to change its action across time slots, and its min-max regret is defined as the difference between its min-max cost and that of the benchmark. The min-max regret is a stringent performance measure and an algorithm with small regret needs to ‘track’ all loss functions simultaneously.
We first show that with adversarial input, min-max regret scales linearly with the time horizon $T$ for any online algorithm. Consequently, we consider a stochastic i.i.d. input model where all loss functions are i.i.d. generated from an unknown joint distribution and propose a simple algorithm that combines the well-known Hedge and online gradient descent (OGD) and show via a remarkably simple proof that its expected min-max regret is $O(\sqrt{T\log(TK)})$. Analogous results are also derived for Martingale difference and Markov input models. The binary prediction problem with multiple binary sequences is also considered where an algorithm has to predict a binary action and the loss function is the maximum of the Hamming distance between the predicted actions and the binary sequences. For this problem, we again show that under adversarial input the min-max regret grows linearly with the time horizon $T$ for any online algorithm, while under the stochastic i.i.d. input model we propose an algorithm that achieves $O(1)$ min-max regret.
@@ 1 Introduction
Online convex optimization (OCO) Hazan (2019) is a basic learning problem that models a wide variety of applications such as spam filtering Hazan (2019), capacity provisioning in cloud networks Lin et al. (2012) etc., where at time $t$, an online algorithm $\mathcal{A}$ has to choose an action $x_{t}\in\mathcal{X}$, after which the adversary reveals the convex objective/loss function $f_{t}$ on which $x_{t}$ is evaluated. As a benchmark, the optimal offline algorithm is considered, that knows the entire set of $f_{t}$, $t=1,\dots,T$, however, can only choose a single action across all times. To quantify the performance of online algorithms with OCO, the metric of static regret $R^{\text{s}ingle}_{\mathcal{A}}(T)=\sup_{f_{t}}\left\{\sum_{t=1}^{T}f_{t}(x_{t})-\min_{x}\sum_{t=1}^{T}f_{t}(x)\right\}$ is studied, where $x_{t}$’s are the actions taken by $\mathcal{A}$.
In a long line of work, comprehensively summarized in Hazan (2019), tight guarantees (both upper and lower bounds) have been obtained for the OCO problem, where optimal $R^{\text{s}ingle}_{\mathcal{A}}(T)=\Theta(\sqrt{T})$. In particular, online gradient descent (OGD), mirror descent, and follow-the-regularized-leader (FTRL) Hazan (2019) all achieve the optimal static regret performance for OCO.
@@
Thus, in this paper, we consider the robust alternative in the multi-objective case: the min-max OCO, where the objective of an algorithm is to minimize the maximum of the total loss (summed across time slots) incurred by each of the multiple objective functions. A compelling example for considering the min-max approach is fair machine learning, where performance is typically evaluated across multiple demographic groups, which naturally induces a multi-objective optimization problem. While Pareto efficiency identifies trade-offs between groups, it may still permit highly unequal solutions. min-max objective effectively addresses this: it maximizes the minimum accuracy (or minimize the maximum error) across groups Chen et al. (2020); Chouldechova and Roth (2020)). See Appendix A for more details on applications of min-max OCO problem.
The formal min-max OCO problem that we consider is as follows. Let there be two convex loss function sequences $f_{t},g_{t},t=1,\dots,T$ that are revealed after $\mathcal{A}$ chooses action $x_{t}$ at time $t$. The min-max cost of $\mathcal{A}$ is
@@
Two loss function sequences are considered for simplicity of exposition since they already capture the basic difficulty. Extension to $K>2$ sequences is straightforward and detailed in Remark 6. As a benchmark, we consider the static offline optimal algorithm $\mathsf{OPT}$ that knows $f_{t},g_{t}$ for $t=1,\dots,T$ at time $t=1$ itself but can only choose a single action $x^{\star}$ across all times and the optimal cost is
$C_{\mathsf{OPT}}=\max\left\{\sum_{t=1}^{T}f_{t}(x^{\star}),\sum_{t=1}^{T}g_{t}(x^{\star})\right\},$ (2)
where $x^{\star}=\arg\min_{x\in\mathcal{X}}\max\left\{\sum_{t=1}^{T}f_{t}(x),\sum_{t=1}^{T}g_{t}(x)\right\}.$ Consequently, we define the min-max static regret of $\mathcal{A}$ with actions $x_{t}$ as
@@
The chosen benchmark in (2) is deliberately stringent and forces the algorithm to track both sequences globally and simultaneously compared to $\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\max\{f_{t}(x),g_{t}(x)\}$, for which the OGD algorithm applied on the surrogate function $h_{t}(x)=\max\{f_{t}(x),g_{t}(x)\}$ would yield regret $O(\sqrt{T})$ independent of the number of distinct loss function sequences.
2 Related Work
Early attempts at multi-objective OCO highlighted the difficulty of defining an appropriate performance metric Mahdavi et al. (2013). A common resolution, introduced in Mahdavi et al. (2013), is constrained online convex optimization (COCO), where one objective is optimized while the remaining objectives are enforced as constraints. In COCO, an algorithm chooses $x_{t}\in\mathcal{X}$ before observing a convex loss $f_{t}$ and convex constraint $g_{t}$, and performance is measured by both static regret with respect to a feasible comparator and cumulative constraint violation. COCO has been extensively studied under i.i.d. inputs Mahdavi et al. (2013); Uziel and El-Yaniv (2017), time-invariant constraints Jenatton et al. (2016); Mahdavi et al. (2012); Yuan and Lamperski (2018); Yi et al. (2021), and fully time-varying constraints using primal–dual or drift-plus-penalty methods Neely and Yu (2017); Guo et al. (2022); Yu et al. (2017), with the best possible guarantee being regret of $O(\sqrt{T})$ and constraint violation of $O(\sqrt{T}\log T)$ with simple algorithms Sinha and Vaze (2024), with further refinements yielding instance-dependent or improved violation bounds Vaze and Sinha (2025); Sinha and Vaze (2025). Dynamic versions of COCO, where the benchmark varies over time, have also been considered Chen and Giannakis (2018); Cao and Liu (2018); Vaze (2022); Liu et al. (2022).
Despite superficial similarities, COCO and min-max OCO are fundamentally different: COCO enforces feasibility over time, whereas min-max OCO compares against a static action that minimizes the worst total loss across objectives. Consequently, results for one setting do not directly imply results for the other. A more detailed discussion on COCO can be found in Appendix B.
@@ 2.1 Global Cost Functions in Experts Setting
A special case of the min-max OCO problem (3) was considered in Even-Dar et al. (2009), where there are $K$ experts, with expert $i$’s loss at time $t$, being $\ell_{t}(i)$. If expert $i$ is selected by an algorithm $\mathcal{A}$ with probability $p_{t}(i)$ at time $t$, then the overall expected loss for expert $i$ is $L_{i}=\sum_{t=1}^{T}p_{t}(i)\ell_{t}(i)$, and the cost of $\mathcal{A}$ is the global cost $C_{\mathcal{A}}=\max_{i=1,\dots,K}\{L_{i}\}$. While Even-Dar et al. (2009) achieved $O(\sqrt{T}\log K)$ regret using the MULTI algorithm, their setting fundamentally differs from ours: their objective is linear in the actions $p_{t}$, and the probability mass $p_{t}(i)$ is split among experts (summing to 1). In contrast, with min-max OCO actions affect all objective functions simultaneously and in full. Consequently, results from Even-Dar et al. (2009) and its extensions Azar et al. (2014); Kesselheim and Singla (2020); Rakhlin et al. (2011) do not apply to our formulation, even for linear functions.
@@ 2.2 Pareto Optimal Multi-Objective OCO
2.3 Per-slot min-max multi-objective optimization
Compared to our ‘global’ min-max objective (2), in Lee et al. (2022), a per-slot min-max multi-objective OCO was studied, where at each round $t=1,\dots,T$: i) an online algorithm chooses $x_{t}\in\mathcal{X}$, ii) the adversary chooses $y_{t}\in\mathcal{Y}$, and iii) the online algorithm suffers a vector loss $\ell_{t}(x_{t},y_{t})=\big(\ell_{t,1}(x_{t},y_{t}),\dots,\ell_{t,d}(x_{t},y_{t})\big)\in[-1,1]^{d}.$ The goal of an online algorithm is to solve $\min_{\{x_{t}\}}\max_{j\in[d]}\sum_{t=1}^{T}\ell_{t,j}(x_{t},y_{t})$, knowing $\ell_{t}(.,.)$, but not $y_{t}$, and its performance is compared against a per-round local minimax value $w_{L}^{t}:=\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\max_{j\in[d]}\ell_{t,j}(x,y),$ where the cumulative benchmark is:
@@ 2.3 Per-slot min-max multi-objective optimization
A Hedge-type algorithm Cesa-Bianchi and Lugosi (2006) was shown in Lee et al. (2022) to have $O(\sqrt{T})$ regret. Defining $\mathcal{Y}=\{e\}$ as a singleton, and $d=2$, such that $\ell_{t,1}(x_{t},y_{t})=f_{t}(x_{t})$ and $\ell_{t,2}(x_{t},y_{t})=g_{t}(x_{t})$ for $y_{t}=e$, we recover our model where two convex functions $f_{t},g_{t}$ are revealed at time $t$, and for which $W_{L}^{T}=\sum_{t=1}^{T}\min_{x_{t}\in\mathcal{X}}\max\{f_{t}(x_{t}),g_{t}(x_{t})\}).$ This benchmark is fundamentally different than our considered benchmark (2) since it is additive across time slots, and allows the benchmark to choose different actions across time slots. Even though this benchmark allows choosing different actions across time slots, surprisingly, we show in Appendix H that $W_{L}^{T}$ can be $3/2$ times more than our benchmark $C_{\mathsf{OPT}}$ (2) and hence algorithms with $O(\sqrt{T})$ regret (5) do not yield a $O(\sqrt{T})$ min-max static regret (3).
2.4 $\alpha$-Fairness
Finally, we review the most closely related problem of online $\alpha$-fairness studied by Si Salem et al. (2023). In this formulation with two users, and at each round $t\in[T]$ an online algorithm selects an action $x_{t}\in\mathcal{X}$. After $x_{t}$ is chosen, two nonnegative concave utility functions $f_{t},g_{t}:\mathcal{X}\to\mathbb{R}_{\geq 0}$ are revealed. Given a sequence of actions $\{x_{t}\}_{t=1}^{T}$, the time-averaged utilities are $\bar{f}(x_{1:T}):=\frac{1}{T}\sum_{t=1}^{T}f_{t}(x_{t})$, and $\bar{g}(x_{1:T}):=\frac{1}{T}\sum_{t=1}^{T}g_{t}(x_{t})$. Let $\bar{f}(x_{1:T})=\bar{f}(x)$ when $x_{t}=x$ for all $t$. Following Si Salem et al. (2023), performance is evaluated via the $\alpha$-fair utility
$U_{\alpha}(\bar{f})+U_{\alpha}(\bar{g}),\qquad U_{\alpha}(z)=\begin{cases}\frac{z^{1-\alpha}}{1-\alpha},&\alpha\neq 1,\\ \log z,&\alpha=1,\end{cases}$ (6)
which is a concave aggregation of the two user utilities. To compare with our min-max framework in this paper, define a cost function $\psi_{\alpha}:\mathbb{R}_{>0}^{2}\to\mathbb{R}$ where $\psi_{\alpha}(a,b):=-\,\Big(U_{\alpha}(a)+U_{\alpha}(b)\Big)$, which is convex in $(a,b)$ for all $\alpha\geq 0$. The offline benchmark of Si Salem et al. (2023) is then $C_{\mathsf{OPT}}(\alpha,T)=\min_{x\in\mathcal{X}}\psi_{\alpha}(\bar{f}(x),\bar{g}(x))$ and the cost of $\mathcal{A}$ is $C_{\mathcal{A}}(\alpha,T)=\psi_{\alpha}(\bar{f}(x_{1:T}),\bar{g}(x_{1:T}))$.
In the limit $\alpha\to\infty$ costs $C_{\mathcal{A}}(\alpha,T)$ and $C_{\mathsf{OPT}}(\alpha,T)$ become equivalent to the min-max costs defined in (1) and (2), respectively. Thus, the online $\alpha$-fairness problem Si Salem et al. (2023) formally subsumes our min-max OCO problem as the limiting case $\alpha\to\infty$.
Si Salem et al. (2023) show that with adversarial input, the regret $C_{\mathcal{A}}(\alpha,T)-C_{\mathsf{OPT}}(\alpha,T)=\Omega(T)$ for any online algorithm, even with fixed $\alpha$, ruling out sublinear regret in this model. Motivated by this impossibility, they consider a stochastic (smoothed) input model, under which they propose an online horizon-fair (OHF) policy and prove that its regret is $O(\alpha\sqrt{T})$. Notably, since the regime of interest for our min-max objective (1) corresponds to $\alpha\to\infty$, the stochastic regret guarantee of Si Salem et al. (2023) becomes vacuous. See Remark 11 for more details.